A 60.0 kg football player leaps straight up into the air (with no horizontal velocity) to catch a pass. He catches the 0.430 kg ball precisely at the peak of his jump, when he is 0.589 meters off the ground. He hits the ground 0.0371 meters away from where he leapt. If the ball was moving horizontally when it was caught, how fast was the ball travelling?
Answer #1

Apply kinmatic equaiton

s = 0.5*a*t^2

0.589= 0.5*9.8*t^2

t = 0.346 s

v*0.346 = 0.0371

v = 0.107m/s

---------------------------------------------------

Apply conservation of momentum

mu = ( mf+ mb) v

0.430 ( u) = (60 + 0.430)0.107

u = 15.03 m/s